Charles Corporation reports the following information for 20X1.

Accounts Receivable – 1/1 $100,000
Accounts Receivable – 12/31 150,000
Inventory – 1/1 40,000
Inventory – 12/31 55,000
Net Credit Sales 800,000
Cost of Goods Sold 450,000
Accumulated Depreciation 15,000

What is the average age of inventory for Charles Corporation?
a. 32.4 days
b. 39.9 days
c. 44.6 days
d. 38.5 days    
e. None of the options listed

please help!thank you!
Answer #1

Average inventory = (40000+55000)/2 = 47500

Inventory turnover ratio = Cost of goods sold/Average inventory = 450000/47500 = 9.47

Average age of inventory = Days in a year/Inventory turnover ratio = 365/9.47 = 38.5 days

Option d is correct.